OpenStudy (anonymous):

Look at the expression. 2/3(x+2)-1/4(x-2) What is an equivalent expression?

OpenStudy (nikato):

\[\frac{ 2 }{ 3(x+2) }-\frac{ 1 }{ 4(x-2) }\]

OpenStudy (nikato):

like this?

OpenStudy (anonymous):

no its set up like 2/3(x+2)-1/4(x-2)

OpenStudy (nikato):

like this then? \[\frac{ 2 }{ 3 }(x+2) - \frac{ 1 }{ 4 }(x-2)\]

OpenStudy (anonymous):

yes

OpenStudy (nikato):

haha. ok then. do you have any answer choices?

OpenStudy (anonymous):

Yes a) 11/12x+22/12 b)11/12x+10/12 c)5/12x+10/12 d)5/12x+22/12

OpenStudy (nikato):

ok. lets look at the first one